OpenStudy (anonymous):

if f(x)= 4+4x-5x^2, find f'(5)

OpenStudy (freckles):

are you allowed to use constant rule,constant multiple rule, power rule?

OpenStudy (freckles):

sum/difference rule?

OpenStudy (freckles):

or do you have to use definition of derivative ?

OpenStudy (anonymous):

yeah i believe so

OpenStudy (freckles):

wait to which one sorry

OpenStudy (freckles):

definition or short cuts?

OpenStudy (anonymous):

not sure, it doesnt say

OpenStudy (freckles):

well if you guys haven't gone over the short cuts then its the definition which so far you have this: \[f'(5)=\lim_{h \rightarrow 0} \frac{f(5+h)-f(5)}{h} \\ f'(5)=\lim_{h \rightarrow 0} \frac{[4-4(5+h)-5(5+h)^2]-[4+4(5)-5(5)^2]}{h}\]

OpenStudy (freckles):

expand and combine like terms

OpenStudy (freckles):

on top

OpenStudy (anonymous):

ok let me see what i get

OpenStudy (anonymous):

ok i got the answer thank you so much @freckles

OpenStudy (freckles):

that was quick

OpenStudy (anonymous):

and i just realized i could of used the short cut, it would have been easier -_-

OpenStudy (freckles):

way easier! :p

OpenStudy (anonymous):

yup xp, oh well @freckles but thnx